# A Survey of 3D Reconstruction with Event Cameras

### v2

@@ III Methods with Stereo Event Cameras
Figure 8: Comparison of depth maps generated from methods using a stereo event camera, adapted from Ghosh et al. [71], under CC BY 4.0, @ Wiley & Sons 2022. (a) SGM method [77] (RGB-based) is adapted to event data by Ghosh et al. [71] (b) GTS method from Leng et al. [36] (c) ESVO method from Zhou et al. [78] (d) Method from Ghosh et al. [71].
III-A Geometry-based Methods
@@
With disparity matching: Many earlier approaches perform depth mapping reconstruction by computing disparity between events observed at the same timestamp across different viewpoints, followed by geometry-based multi-view stereo estimation to achieve real-time 3D depth reconstruction. In 2015, Schraml et al. [32] introduced a novel stereo matching method for a rotating panoramic stereo event system. Their approach defines a cost metric based on event distribution similarity between left and right views and computes disparity through dynamic programming on sparse event maps. The resulting disparities are then used to reconstruct real-world depth in 360° panoramic views. In 2018, Zhu et al. [37] proposed a method that synchronises events in time using known camera velocities, constructs a dense event disparity volume, and performs real-time sliding window matching, introducing a novel matching cost function combining ambiguity and similarity. Simultaneously, Leng et al. [36] reformulated event matching as a time-based stereo vision problem, where disparity estimation is achieved by detecting the temporal coincidence of events across epipolar lines. Their method introduces a generalised spiking neuron model that integrates spatio-temporal event information with temporal decay, allowing robust matching under varying motion and illumination conditions. A voting scheme is then used to infer depth from the accumulated neural activations, enabling dense 3D reconstruction from pure event streams. In 2025, Freitag et al. [79] proposed a stereo reconstruction method that performs disparity matching by leveraging temporally coinciding events between two event cameras. By triangulating matched event pairs, their system reconstructs high-precision point clouds.
TABLE II: Methods with stereo event cameras
Author Year Priors Event Rep. Real-time Output (Dense?) Dataset
Schraml et al. [32] 2015 Trajectory EAF 3D depth map (✗) Self-collected [32]
Zhu et al. [37] 2018 Velocity Event disparity volume 3D depth map (✓) MVSEC [80]
Zhou et al. [81] 2018 Pose Time surface 3D depth map (✗) MVSEC [80]
Leng et al. [36] 2018 Pose Time surface Point cloud, depth map (✗) Self-collected [36]
Zhu et al. [82] 2019 Pose Event disparity volume 3D depth map (✗) Self-collected [72]
Steffen et al. [72] 2019 - 4D vector Voxel (✗) MVSEC [80]
Zhou et al. [78] 2021 Pose Time surface Point cloud, depth map (✗) MVSEC [80]
Nam et al. [73] 2022 Pose Event Stack 3D depth map (✓) DSEC [83]
Ghosh et al. [71] 2022 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83], etc
Ghosh et al. [84] 2024 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83], etc
Freitag et al. [79] 2025 Pose Event Stack Point cloud (✓) Self-collected [85]
Hitzges et al. [86] 2025 Pose Event disparity volume 3D depth map (✗) MVSEC [80], DSEC [83]
@@
Without disparity matching: Disparity matching can also be bypassed altogether by directly optimising geometric or temporal consistency across event streams. In 2018, Zhou et al. [81] proposed a novel forward-projection-based depth estimation method, which avoids the need for explicit disparity computation. Instead of finding correspondences between stereo events, their method optimises a temporal consistency energy by projecting candidate depth hypotheses into both stereo time surfaces and measuring their alignment. A coarse-to-fine search strategy is used to obtain semi-dense depth maps in real-time, enabling robust depth reconstruction under challenging conditions. Building upon this idea, Zhou et al. [78] extended the concept to a stereo event-based visual odometry framework. By jointly optimising over spatial and temporal consistency across stereo time surfaces, they estimate not only per-frame depth but also the stereo camera motion. The system runs in real-time and outputs semi-dense depth maps aligned with visual odometry estimates, making it well-suited for 3D reconstruction in dynamic environments. Later in 2022, Ghosh et al. [71] proposed a multi-event-camera fusion framework that leverages Disparity Space Images (DSIs) [87] to accumulate event ray densities from multiple viewpoints. Instead of requiring event correspondences, their method refocuses events in DSI space and performs probabilistic fusion to obtain high-quality depth maps. They further introduce an outlier rejection strategy to enhance robustness and demonstrate generalisation across multiple datasets without fine-tuning. Building upon this idea, in 2024, Ghosh et al. [84] proposed ES-PTAM, which combines a mapping module based on ray density fusion and a tracking module using edge-map alignment, both operating on pure event data. By bypassing disparity matching and processing stereo events in a correspondence-free manner via geometric ray density fusion, the system produces accurate semi-dense depth maps and camera poses in real time. It is a leading stereo visual odometry method that exclusively uses events, without any auxiliary inputs.
III-B Learning-based Methods
@@
Unsupervised structural embedding: Additionally, one research explores unsupervised clustering for self-organising structural modelling. In 2019, Steffen et al. [72] proposed an unsupervised reconstruction framework using Self-Organising Maps (SOMs) [88]. By embedding stereo events from multiple viewpoints into a shared latent space, their method performs sparse voxel-based 3D reconstruction without requiring calibration or supervision, offering a lightweight and biologically inspired approach to structural modelling.
Deep learning processing: More recent methods leverage deep neural networks to learn stereo disparity estimation and event representations. In 2019, Zhu et al. [82] proposed an unsupervised deep learning framework, which leverages neural networks to predict depth and motion from stereo event data. They introduce a discretized event volume representation to preserve spatio-temporal structure and apply a motion compensation loss that reduces event blur. In 2022, Nam et al. [73] introduced a deep stereo framework that combines multi-density event stacking with attention-guided encoding via a UNet-ResNet [89, 90] backbone. Their method further incorporates future event prediction during training through a distillation loss [91, 92], enabling real-time, dense depth estimation with high accuracy on the DSEC dataset [93]. In 2025, Hitzges et al. [86] proposed DERD-Net, which estimates depth from event-based ray densities by processing DSIs derived from multi-view event data and known camera poses. By extracting local sub-volumes (Sub-DSIs) and combining 3D convolutions with recurrent units, their method enables efficient, high-resolution depth prediction.
IV Methods with Monocular Event Cameras
Figure 9: Overview of methods with monocular event camera that generate traditional output. (a) Schematic of spatial scanning in a 3D scene (b) Schematic of static event camera scanning rotated object on a platform (c) The E3D pipeline that uses E2S to estimate physical prior - Silhouette, to assist 3D mesh reconstruction [94] (d) The end-to-end structure that directly generate 3D model from Chen et al. [95]
@@ IV Methods with Monocular Event Cameras
Existing monocular approaches can be broadly divided into four categories based on their underlying modelling techniques and output formats: (IV-A) geometry-based methods with traditional outputs, (IV-B) learning-based methods with traditional outputs, (IV-C) methods with Neural Radiance Fields, and (IV-D) methods with 3D Gaussian Splatting. The first two categories aim to reconstruct explicit 3D outputs such as depth maps, point clouds, meshes, or voxels, while the latter two leverage volumetric rendering paradigms to generate photorealistic and continuous scene representations from asynchronous event streams.
Additionally, some methods leverage RGB Bayer event streams captured by colour event sensors like Colour DAVIS (e.g., DAVIS 346c) to support colour-consistent NeRF and Gaussian rendering [96, 97, 98]. However, since the colour is encoded within the same event modality, these approaches are still considered monocular, not multimodal.
IV-A Geometry-based Methods with Traditional Output
TABLE III: Monocular event camera: Geometry-based methods with traditional outputs
@@ IV-A Geometry-based Methods with Traditional Output
Kim et al. [33] 2016 Trajectory EGA 3D depth map (✗) Self-collected [33]
@@
Rebecq et al. [99] 2016 Trajectory Event-by-event 3D depth map (✗) Self-collected [51]
Rebecq et al. [100] 2016 Pose Event frame 3D depth map (✗) Self-collected [100]
Guan et al. [28] 2024 Trajectory Time surface 3D depth map (✓) DAVIS240C [101]
Elms et al. [102] 2024 Trajectory Time surface Sparse point cloud (✗) TOPSPIN [102]
@@
TABLE IV: Monocular event camera: Learning-based methods with traditional outputs
@@
Author Year Priors Pipeline Event Rep. Model Output Dataset
Baudron et al. [94] 2020 Silhouette EAF E2S(CNN) Mesh (✓) ShapeNet [103]
Xiao et al. [104] 2022 Pose Voxel grid E2VID(RNN-CNN) Mesh (✓) ESIM [105]
@@
Wang et al. [106] 2023 Contour, Trajec. Voxel grid Evac3d(CNN) Mesh (✓) MOEC-3D [106]
Chen et al. [95] 2023 - Event frame E2V(CNN) Voxel (✓) SynthEVox3D [95]
Xu et al. [27] 2025 - Event frame (sobel) E2V(CNN) Voxel (✓) SynthEVox3D [95]
Geometry-based methods typically achieve semi-dense or sparse 3D reconstruction in real-time, relying on spatial-temporal event aggregation and geometric priors. Since monocular event cameras cannot directly observe disparities, these methods often require known or estimated motion to compensate for the view angles. Figure 9 and Table III provide an overview of these methods. Based on the data acquisition strategy, we categorise these methods into the following two types:
Spatial scanning: Many methods rely on spatial scanning, where the event camera must move through space to accumulate multi-view observations. By capturing asynchronous events from different viewpoints, these methods infer scene structure through motion-induced parallax. In 2016, Kim et al. [33] proposed an approach utilising three interleaved probabilistic filters to estimate camera trajectory, scene log-intensity gradient, and inverse depth. In 2016, Rebecq et al. [99] proposed EMVS, employing event space-sweep and ray density analysis to directly generate a semi-dense 3D depth map, without frame-level data association. Later, EMVS was extended into a full SLAM framework as its mapping module, where Rebecq et al. [100] introduced EVO (Event-based Visual Odometry), which integrates event projection, time-surface alignment, and DSI-based depth fusion to jointly estimate camera trajectory and reconstruct 3D geometry in real time. A recent study has improved geometry-based methods to achieve dense reconstruction. In 2024, Guan et al. [28] proposed EVI-SAM, a tightly coupled event-image-IMU SLAM framework. It achieves real-time dense 3D reconstruction on a standard CPU, integrating event-based 2D-2D alignment, image-guided depth interpolation, and TSDF fusion [107].
Static camera scanning rotated object: A unique innovation enables reconstruction when the event camera is stationary, while the object rotates. In 2024, Elms et al. [102] proposed eSfO, which performs 3D reconstruction through event corner tracking and factor graph optimisation [108], but it only performs non-real-time sparse point cloud reconstruction.
IV-B Learning-Based Methods with Traditional Output
@@
Deep learning-based methods typically produce non-real-time dense reconstruction. However, traditional RGB image feature extraction techniques cannot be directly applied to raw event data, and image-based event representation is often required as a preprocessing [109]. Table IV and Figure 9 provide an overview of these methods. Based on the reliance on structured pipelines and external priors, we categorise these methods into the following two types:
@@
With pipeline and priors: Many studies have established an event-to-3D pipeline, a structured and modular event processing framework [94, 104, 109, 100], including feature extraction, matching, and 3D computation. The extraction and estimation of priors are also essential. In 2020, Baudron et al. [94] proposed E3D, the first dense 3D shape reconstruction method based on monocular event cameras. It employs the E2S neural network to estimate silhouettes and leverages PyTorch3D [110] for 3D mesh optimisation, achieving high-quality multi-view 3D reconstruction trained on ShapeNet [111]. In 2022, Xiao et al. [104] proposed a pipeline using the E2VID deep learning method [112] to process continuous event streams and generate normalised intensity image sequences. They then employed SfM [113] to estimate intrinsic and extrinsic parameters for sparse point clouds and used MVS [114, 115, 116, 117] for dense mesh reconstruction. In 2023, Wang et al. [106] proposed EvAC3D, which uses CNN to predict Apparent Contour Events (ACE), combined with Continuous Volume Carving and Global Mesh Optimisation, to achieve dense 3D shape reconstruction with known camera trajectories.

### v3

@@ 3 Methods with Stereo Event Cameras
Figure 8: Comparison of depth maps generated from methods using a stereo event camera, adapted from Ghosh et al., under $C$ BY 4.0, @ Wiley & Sons 2022. (a) SGM method (RGB-based) is adapted to event data by Ghosh et al. (b) GTS method from Leng et al. (c) ESVO method from Zhou et al. (d) Method from Ghosh et al.
Table 2: Methods with stereo event cameras
Author Year Priors Event Rep. Real-time Output (Dense?) Dataset
Schraml et al. [[S9]] 2015 Trajectory EAF 3D depth map (✗) Self-collected [[S9]]
Zhu et al. [[S0]] 2018 Velocity Event disparity volume 3D depth map (✓) MVSEC [MVSEC]
Zhou et al. [[S2]] 2018 Pose Time surface 3D depth map (✗) MVSEC [MVSEC]
Leng et al. [[S4]] 2018 Pose Time surface Point cloud (✗) Self-collected [[S4]]
Zhu et al. [[XS2]] 2019 Pose Event disparity volume 3D depth map (✗) Self-collected [[S5]]
Steffen et al. [[S5]] 2019 - 4D vector Voxel (✗) MVSEC [MVSEC]
Zhou et al. [[S6]] 2021 Pose Time surface Point cloud (✗) MVSEC [MVSEC]
Nam et al. [[S7]] 2022 Pose Event Stack 3D depth map (✓) DSEC [DSEC2]
Ghosh et al. [[S8]] 2022 Pose Event disparity volume 3D depth map (✗) MVSEC [MVSEC], DSEC [DSEC2], etc
Ghosh et al. [[XS3]] 2024 Pose Event disparity volume 3D depth map (✗) MVSEC [MVSEC], DSEC [DSEC2], etc
Freitag et al. [[XS1]] 2025 Pose Event Stack Point cloud (✓) Self-collected [[XS1-data]]
Hitzges et al. [[XS4]] 2025 Pose Event disparity volume 3D depth map (✗) MVSEC [MVSEC], DSEC [DSEC2]
3.1 Geometry-based Methods
@@ 3.1.1 With disparity matching
Many earlier approaches perform depth mapping reconstruction by computing disparity between events observed at the same timestamp across different viewpoints, followed by geometry-based multi-view stereo estimation to achieve real-time 3D depth reconstruction. In 2015, Schraml et al. [[S9]] introduced a novel stereo matching method for a rotating panoramic stereo event system. Their approach defines a cost metric based on event distribution similarity between left and right views and computes disparity through dynamic programming on sparse event maps. The resulting disparities are then used to reconstruct real-world depth in 360° panoramic views. In 2018, Zhu et al. [[S0]] proposed a method that synchronises events in time using known camera velocities, constructs a dense event disparity volume, and performs real-time sliding window matching, introducing a novel matching cost function combining ambiguity and similarity. Simultaneously, Leng et al. [[S4]] reformulated event matching as a time-based stereo vision problem, where disparity estimation is achieved by detecting the temporal coincidence of events across epipolar lines. Their method introduces a generalised spiking neuron model that integrates spatio-temporal event information with temporal decay, allowing robust matching under varying motion and illumination conditions. A voting scheme is then used to infer depth from the accumulated neural activations, enabling dense 3D reconstruction from pure event streams. In 2025, Freitag et al. [[XS1]] proposed a stereo reconstruction method that performs disparity matching by leveraging temporally coinciding events between two event cameras. By triangulating matched event pairs, their system reconstructs high-precision point clouds.
3.1.2 Without disparity matching
@@
Disparity matching can also be bypassed altogether by directly optimising geometric or temporal consistency across event streams. In 2018, Zhou et al. [[S2]] proposed a novel forward-projection-based depth estimation method, which avoids the need for explicit disparity computation. Instead of finding correspondences between stereo events, their method optimises a temporal consistency energy by projecting candidate depth hypotheses into both stereo time surfaces and measuring their alignment. A coarse-to-fine search strategy is used to obtain semi-dense depth maps in real-time, enabling robust depth reconstruction under challenging conditions. Building upon this idea, Zhou et al. [[S6]] extended the concept to a stereo event-based visual odometry framework. By jointly optimising over spatial and temporal consistency across stereo time surfaces, they estimate not only per-frame depth but also the stereo camera motion. The system runs in real-time and outputs semi-dense depth maps aligned with visual odometry estimates, making it well-suited for 3D reconstruction in dynamic environments. Later in 2022, Ghosh et al. [[S8]] proposed a multi-event-camera fusion framework that leverages Disparity Space Images (DSIs) [DSI] to accumulate event ray densities from multiple viewpoints. Instead of requiring event correspondences, their method refocuses events in DSI space and performs probabilistic fusion to obtain high-quality depth maps. They further introduce an outlier rejection strategy to enhance robustness and demonstrate generalisation across multiple datasets without fine-tuning. Building upon this idea, in 2024, Ghosh et al. [[XS3]] proposed ES-PTAM, which combines a mapping module based on ray density fusion and a tracking module using edge-map alignment, both operating on pure event data. By bypassing disparity matching and processing stereo events in a correspondence-free manner via geometric ray density fusion, the system produces accurate semi-dense depth maps and camera poses in real time. It is a leading stereo visual odometry method that exclusively uses events, without any auxiliary inputs.
3.2 Learning-based Methods
@@
3.2.1 Unsupervised structural embedding
Additionally, one research explores unsupervised clustering for self-organising structural modelling. In 2019, Steffen et al. [[S5]] proposed an unsupervised reconstruction framework using Self-Organising Maps (SOMs) [[S5-13]]. By embedding stereo events from multiple viewpoints into a shared latent space, their method performs sparse voxel-based 3D reconstruction without requiring calibration or supervision, offering a lightweight and biologically inspired approach to structural modelling.
3.2.2 Deep learning processing
More recent methods leverage deep neural networks to learn stereo disparity estimation and event representations. In 2019, Zhu et al. [[XS2]] proposed an unsupervised deep learning framework, which leverages neural networks to predict depth and motion from stereo event data. They introduce a discretized event volume representation to preserve spatio-temporal structure and apply a motion compensation loss that reduces event blur. In 2022, Nam et al. [[S7]] introduced a deep stereo framework that combines multi-density event stacking with attention-guided encoding via a UNet-ResNet [[S7-35], [S7-14]] backbone. Their method further incorporates future event prediction during training through a distillation loss [[S7-15], [S7-38]], enabling real-time, dense depth estimation with high accuracy on the DSEC dataset [[S7-13]]. In 2025, Hitzges et al. [[XS4]] proposed DERD-Net, which estimates depth from event-based ray densities by processing DSIs derived from multi-view event data and known camera poses. By extracting local sub-volumes (Sub-DSIs) and combining 3D convolutions with recurrent units, their method enables efficient, high-resolution depth prediction.
Figure 9: Overview of methods with monocular event camera that generate traditional output. (a) Schematic of spatial scanning in a 3D scene. (b) Schematic of static event camera scanning rotated object on a platform. (c) The E3D pipeline that uses E2S to estimate physical prior - Silhouette, to assist 3D mesh reconstruction. (d) The end-to-end structure that directly generate 3D model from Chen et al.
4 Methods with Monocular Event Cameras
@@
Existing monocular approaches can be broadly divided into four categories based on their underlying modelling techniques and output formats: (4.1) geometry-based methods with traditional outputs, (4.2) learning-based methods with traditional outputs, (4.3) methods with Neural Radiance Fields, and (4.4) methods with 3D Gaussian Splatting. The first two categories aim to reconstruct explicit 3D outputs such as depth maps, point clouds, meshes, or voxels, while the latter two leverage volumetric rendering paradigms to generate photorealistic and continuous scene representations from asynchronous event streams.
Additionally, some methods leverage RGB Bayer event streams captured by colour event sensors like Colour DAVIS (e.g., DAVIS 346c) to support colour-consistent NeRF and Gaussian rendering [[N2], [G5], [G11]]. However, since the colour is encoded within the same event modality, these approaches are still considered monocular, not multimodal.
4.1 Geometry-based Methods with Traditional Output
Geometry-based methods typically achieve semi-dense or sparse 3D reconstruction in real-time, relying on spatial-temporal event aggregation and geometric priors. Since monocular event cameras cannot directly observe disparities, these methods often require known or estimated motion to compensate for the view angles. Figure 9 and Table 3 provide an overview of these methods. Based on the data acquisition strategy, we categorise these methods into the following two types:
4.1.1 Spatial scanning
Many methods rely on spatial scanning, where the event camera must move through space to accumulate multi-view observations. By capturing asynchronous events from different viewpoints, these methods infer scene structure through motion-induced parallax. In 2016, Kim et al. [[M1]] proposed an approach utilising three interleaved probabilistic filters to estimate camera trajectory, scene log-intensity gradient, and inverse depth. In 2016, Rebecq et al. [[EMVSconf]] proposed EMVS, employing event space-sweep and ray density analysis to directly generate a semi-dense 3D depth map, without frame-level data association. Later, EMVS was extended into a full SLAM framework as its mapping module, where Rebecq et al. [[M2]] introduced EVO (Event-based Visual Odometry), which integrates event projection, time-surface alignment, and DSI-based depth fusion to jointly estimate camera trajectory and reconstruct 3D geometry in real time. A recent study has improved geometry-based methods to achieve dense reconstruction. In 2024, Guan et al. [[M3]] proposed EVI-SAM, a tightly coupled event-image-IMU SLAM framework. It achieves real-time dense 3D reconstruction on a standard CPU, integrating event-based 2D-2D alignment, image-guided depth interpolation, and TSDF fusion [[M3-58]].
4.1.2 Static camera scanning rotated object
A unique innovation enables reconstruction when the event camera is stationary, while the object rotates. In 2024, Elms et al. [[M4]] proposed eSfO, which performs 3D reconstruction through event corner tracking and factor graph optimisation [[M4-12]], but it only performs non-real-time sparse point cloud reconstruction.
Table 3: Monocular event camera: Geometry-based methods with traditional outputs
@@ 4.1.2 Static camera scanning rotated object
Kim et al. [[M1]] 2016 Trajectory EGA 3D depth map (✗) Self-collected [[M1]]
@@
Rebecq et al. [[EMVSconf]] 2016 Trajectory Event-by-event 3D depth map (✗) Self-collected [[EMVS]]
Rebecq et al. [[M2]] 2016 Pose Event frame 3D depth map (✗) Self-collected [[M2]]
Guan et al. [[M3]] 2024 Trajectory Time surface 3D depth map (✓) DAVIS240C [DAVIS240C]
Elms et al. [[M4]] 2024 Trajectory Time surface Point cloud (✗) TOPSPIN [[M4]]
@@
Table 4: Monocular event camera: Learning-based methods with traditional outputs
@@
Author Year Priors Pipeline Event Rep. Model Output (Dense?) Dataset
Baudron et al. [[MD1]] 2020 Silhouette EAF E2S(CNN) Mesh (✓) ShapeNet [[MD1-D]]
Xiao et al. [[MD2]] 2022 Pose Voxel grid E2VID(RNN-CNN) Mesh (✓) ESIM [[esim]]
@@
Wang et al. [[MD5]] 2023 Contour, Trajec. Voxel grid Evac3d(CNN) Mesh (✓) MOEC-3D [[MD5]]
Chen et al. [[Dense]] 2023 - Event frame E2V(CNN) Voxel (✓) SynthEVox3D [[Dense]]
Xu et al. [[Towards]] 2025 - Event frame E2V(CNN) Voxel (✓) SynthEVox3D [[Dense]]
4.2 Learning-Based Methods with Traditional Output
@@
Deep learning-based methods typically produce non-real-time dense reconstruction. However, traditional RGB image feature extraction techniques cannot be directly applied to raw event data, and image-based event representation is often required as a preprocessing [[N7]]. Table 4 and Figure 9 provide an overview of these methods. Based on the reliance on structured pipelines and external priors, we categorise these methods into the following two types:
4.2.1 With pipeline and priors
@@
Many studies have established an event-to-3D pipeline, a structured and modular event processing framework [[MD1], [MD2], [N7], [M2]], including feature extraction, matching, and 3D computation. The extraction and estimation of priors are also essential. In 2020, Baudron et al. [[MD1]] proposed E3D, the first dense 3D shape reconstruction method based on monocular event cameras. It employs the E2S neural network to estimate silhouettes and leverages PyTorch3D [[MD1-29]] for 3D mesh optimisation, achieving high-quality multi-view 3D reconstruction trained on ShapeNet [[MD1-5]]. In 2022, Xiao et al. [[MD2]] proposed a pipeline using the E2VID deep learning method [[T4]] to process continuous event streams and generate normalised intensity image sequences. They then employed SfM [[MD2-17]] to estimate intrinsic and extrinsic parameters for sparse point clouds and used MVS [[MD2-28], [MD2-29], [MD2-30], [MD2-31]] for dense mesh reconstruction. In 2023, Wang et al. [[MD5]] proposed EvAC3D, which uses CNN to predict Apparent Contour Events (ACE), combined with Continuous Volume Carving and Global Mesh Optimisation, to achieve dense 3D shape reconstruction with known camera trajectories.
4.2.2 Without pipeline and priors